OpenStudy (anonymous):

What is the simplest form of the expression below 2x^2-10x-28 x 6 ------------ ----- 6x x-7

OpenStudy (anonymous):

the 6 in the right numerator cancels the one in the left denominator

OpenStudy (anonymous):

so you're left with x(x-7) in the denominator, right?

OpenStudy (anonymous):

now can you factorize the numerator 2x^2-10x-28 = ?

OpenStudy (anonymous):

ya 2x^2-10x-28 ------------ x-7 right?

OpenStudy (anonymous):

no only the 6 's cancel each other, the X remains, so the denominator will become x(x-7)

OpenStudy (anonymous):

oh okay so 2 (x+2) (x-7)

OpenStudy (anonymous):

@kx2bay

OpenStudy (anonymous):

huh? sorry im bad at math

OpenStudy (anonymous):

yes, correct

OpenStudy (anonymous):

okay but that not the answer @kx2bay

OpenStudy (anonymous):

right? i have t simplify

OpenStudy (anonymous):

so what's left...? you're not that bad! you managed to factorize the main part

OpenStudy (anonymous):

\[2(x+2)(x-7) and (x-7)\]

OpenStudy (anonymous):

you have (x-7) in the numerator and in the denominator, don't they cancel each other out? if yes, then what remains in the numerator and the denominator?

OpenStudy (anonymous):

2(x+2)

OpenStudy (anonymous):

which i would mutiply right?

OpenStudy (anonymous):

you keep forgetting the x in the denominator, why? it was part of 6x initially and only the 6 was cancelled out

OpenStudy (anonymous):

oh ya sorry because im not writing on peice of paper i do online classes

OpenStudy (anonymous):

2(x+2) ------ x

OpenStudy (anonymous):

always use a piece of paper ;)

OpenStudy (anonymous):

that's correct, well done!!! Champ :)

OpenStudy (anonymous):

so would it be 2x+4 ----- x @kx2bay

OpenStudy (anonymous):

either that or 2(x+2) ------ x I'm not sure which is simpler, to me both are simple ;)

OpenStudy (anonymous):

can you help some more? @kx2bay

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\frac{ x^2yz }{ y^2 } \times \frac{ y }{ 2x}\]

OpenStudy (anonymous):

@kx2bay

OpenStudy (anonymous):

y from right numer. cancels one from the left denom., what's left?

OpenStudy (anonymous):

x^2yz ------- 2xy @kx2bay

OpenStudy (anonymous):

beauty, now can't u do the same with the x's ?

OpenStudy (anonymous):

xyz ------ 2y

OpenStudy (anonymous):

well done, what about the y's now?

OpenStudy (anonymous):

xz ______ 2y^2

OpenStudy (anonymous):

but why does the the other y go to the bottom and what are these problems called factoring?

OpenStudy (anonymous):

why did you put y squared in the denom.? I thought the y's cancelled each other out :)

OpenStudy (anonymous):

factoring = breaking up a big expression into smaller expressions called factors

OpenStudy (anonymous):

so xz ---- 2

OpenStudy (anonymous):

Yes that's PERFECT I hope I didn't confuse u above

OpenStudy (anonymous):

f^2-25 x f ------ ---- 5 f-5 @kx2bay

OpenStudy (anonymous):

start small :)

OpenStudy (anonymous):

do you know how to factorize f^2 - 25 25 can be represented as 5^2

OpenStudy (anonymous):

(f-5) (f-5)

OpenStudy (anonymous):

this is called difference between 2 sqaures a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

you almost got it just got the wrong sign ;)

OpenStudy (anonymous):

(f+5)(f+5)

OpenStudy (anonymous):

no buddy, one has a - and the other +

OpenStudy (anonymous):

look at this: a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

ahh okay (f-5)(f+5) f^2-5 ------- 5

OpenStudy (anonymous):

was that the final answer that u wrote above, bit squashed next the f+5 ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

no, don't rush, take a step back, use a pen and paper it will help

OpenStudy (anonymous):

(f-5)(f+5) f ------- . -------- 5 (f-5)

OpenStudy (anonymous):

(f-5) cancel each other from the numer. and denom. correct ? if yes then can u write what's left?

OpenStudy (anonymous):

f/5

OpenStudy (anonymous):

and where did the (f+5) go? vanished :)

OpenStudy (anonymous):

the answer is ...

OpenStudy (anonymous):

f^2+5/5

OpenStudy (anonymous):

go back a step and put it in the form of (f+5).f/5

OpenStudy (anonymous):

now f * (f+5) = f^2 + 5f do you agree?

OpenStudy (anonymous):

ya soo F^2+5f ---------

OpenStudy (anonymous):

5

OpenStudy (anonymous):

what are these problems called

OpenStudy (anonymous):

the answer is either f(f+5) ------ 5 or f^2 + 5f --------- 5

OpenStudy (anonymous):

factorizing and simplifying

OpenStudy (anonymous):

good luck @Claytonwj4 me go sleep now